# Neural Networks for Extreme Quantile Regression with an Application to Forecasting of Flood Risk

### v3

@@ 4.2 Results
For estimation of the conditional GPD parameters with our recurrent EQRN model, we consider both LSTM and GRU architectures with one to three recurrent layers and hidden dimensions between $32$ and $256$. As the networks are not too deep, $L_{2}$ penalty was chosen over dropout for regularization during training, with possible penalty $\lambda\in\{0,10^{-6},10^{-5},10^{-4},10^{-3}\}$. Both constant and covariate-dependent shape parameter outputs are considered. The model with minimum validation loss is a single LSTM layer with hidden dimension $128$, followed by the usual fully connected output layer, with constant shape and $\lambda=10^{-4}$. As a comparison, we also retain predictions for the best unpenalized network with $\lambda=0$ and fixed shape, which has two LSTM layers of hidden size $128$.
The left panel of Figure 5 shows the RMSE of best penalized and unpenalized EQRN models compared to the improved competitors as a function of the quantile level $\tau$. We observe that for the lowest level $\tau=\tau_{0}=80\%$, all structured GPD models have the same performance since they use the same intermediate quantiles. GRF and the unconditional model have already higher errors since they are not able to capture the sequential dependence at the intermediate level sufficiently. For growing quantile levels $\tau$, the errors of the covariate-dependent GPD models start to diverge. This is due to the differences in modelling flexibility in terms of the GPD parameters of each method. We observe a similar behaviour for EXQAR, as its linear quantile dependence is not flexible enough. Our EQRN method based on recurrent neural networks seems to be best at modelling sequential tail dependence.
Figure 5: Left: root mean squared error between predicted and true conditional quantiles at different probability levels $\tau$ (log-scale), for the selected EQRN models and the improved competitors. Centre-right: true versus predicted quantiles at probability level $\tau=99.95\%$ for the best unpenalized (middle) and penalized (right) EQRN models (dots), compared to the semi-conditional estimates (crosses).
Figure 5 also shows the predicted quantiles $\hat{Q}_{\tilde{\mathbf{x}}_{u}}(\tau)$ on the test data compared to the true $Q_{\tilde{\mathbf{x}}_{u}}(\tau)$ for a fixed $\tau=99.95\%$ for the best penalized and unpenalized EQRN models. In general, both models seem to perform well in predicting the high conditional quantiles. The weight penalty seems to mainly affect the larger quantile predictions. Compared to the unpenalized model, we observe that the reduction in the variance of the predictions comes at the cost of a bias for larger quantile values. This bias-variance trade-off is typical with penalization. The poor performance of the semi-conditional estimates highlights the added value of covariate dependence in the GPD parameters.
@@ 5.1 Motivation
Flood risk is a major natural hazard in Europe, which causes huge economic damage and endangers human lives. There is a longstanding interest in statistical methods of extreme value theory for hydrology (e.g., Katz et al., 2002; Keef et al., 2009; Asadi et al., 2015; Engelke and Hitz, 2020), and national agencies commonly use them to assess the long-term risk of flooding in cities, at power plants and other key locations. Return levels with long return periods can be estimated using the GEV distribution for annual maxima or the GPD model for daily threshold exceedances. The output then guides effective long-term flood management measures.
An example of an important location in Switzerland is the gauging station in Bern on the Aare river, which is shown within its water catchment in Figure 1. The Swiss Federal Office for the Environment (FOEN) monitors the Aare, and we use daily average discharges (in $\text{m}^{3}\text{s}^{-1}$) in Bern and another upstream station together with recordings of daily precipitation (in mm) at six locations in the Bern catchment; see Figure 1 for details. All time series are available in the period from 1930–2014 and can be obtained from the FOEN [^2] https://www.hydrodaten.admin.ch/ (for discharges) and MeteoSwiss [^3] https://gate.meteoswiss.ch/idaweb (for precipitation). Figure 6 shows an excerpt for the two river stations and one precipitation gauge.
To illustrate possible drawbacks of a classical extreme value analysis, the left panel of Figure 7 shows the annual maxima of river discharges at the Bernese station on the Aare. The dashed line is the estimated 100-year return level based on the GEV approximation using the training period from 1930–1958. One can see that starting from the year 1999 there are several exceedances over this return level, somewhat contradicting the fact that it should only be exceeded on average once in 100 years. The solid line is the same return level based on data from 1930–$y$, where $y\in\{1959,\dots,2014\}$ denotes the end of the training period. While the predictions are fairly stable until 1999, an extreme value analysis performed after that year would yield much higher values for the 100-year return level. Conversely, historical estimates before such a break-point would severely underestimate the flood risk. In general, distributional shifts can be due to climate change, changes in the river system or other structural breaks in factors influencing discharge at this location. For the Bernese station, the FOEN indeed reports a significant break-point in extreme discharges in the nineties but acknowledges that a clear cause can not be identified444see flood report of the FOEN at https://www.hydrodaten.admin.ch/en/2135.html. One factor may be a multi-decadal variability of flood occurrence, as described in Schmocker-Fackel and Naef (2010).
@@
Figure 7: Left: annual maxima of daily average discharges (points) at Bern–Schönau (62) together with the unconditional 100-year return level based on GEV fitted on the training data of 1930–1958 (dashed line), and the evolution of the same return level (solid line) using data from 1930–$y$, where $y\in\{1959,\dots,2014\}$ denotes end of the period. Middle: evolution of the validation loss (solid line) of the selected EQRN network for the river discharge data as a function of the training epoch; the dashed line shows the validation loss of the semi-conditional model. Right: number of observations exceeding the EQRN quantile predictions on the test set (dotted line) compared to the expected number of exceedances (solid line) for different probability levels (log-scale).
The aim of our methodology is complementary to classical extreme value analysis and addresses this issue with static return levels. We apply our EQRN model to estimate one-day-ahead extreme quantiles of the river discharge conditionally on previous observations of discharge and precipitation in the catchment. This allows the forecasting of flood risk even in non-stationary systems, such as a changing climate. The strength of our approach lies in the ability to exploit information from multiple covariates and capture the complex time dependence. Even in situations where the causes for structural changes are unknown, our method implicitly accounts for them through their effects on the covariates. The output of the model can help practitioners and authorities to manage flooding more effectively and help to minimize their disastrous impacts by early warning systems.
@@ 5.2 Model Specification
The whole dataset consists of 31,046 daily observations $(\mathbf{x}_{t},y_{t})$ between 1930–2014. The response $y_{t}$ is the daily average discharge at the Bernese gauging station on the Aare, and the covariates $\mathbf{x}_{t}\in\mathbb{R}^{p}$, $p=7$, consist of discharge at another upstream station and daily precipitation measurements from six locations in the same catchment; see Figures 1 and 6. The discharges show significant seasonality, both in trend and variance, with the largest extremes only appearing in the summer. We do not reduce artificially the non-stationarity via classical approaches from times series analysis (e.g., Cleveland et al., 1990), as we believe the seasonality and other trends are captured through the covariates.
We split the data into training and test sets. The first $T=10,349$ observations in the period between 1930–1958 are used to train the models, whereof the first three-quarters serve the parameter estimation, and the remaining quarter is a validation set to determine hyperparameters (sets $\mathcal{T}$ and $\mathcal{V}$ in Algorithm 2, respectively). The test set contains 20,697 observations from 1958–2014, which is used for neither fitting nor selection of parameters, but only to evaluate the model performance on an independent time period. We choose this rather small proportion of training data to study the ability of the model to adapt to possible non-stationarity over time without refitting. In particular, the model weights are not updated with any information from data after 1958, even for forecasts in the study of the 2005 flood of interest. A large test set is also required to evaluate extreme properties of the data distribution. As augmented covariates at time $t$ for the recurrent neural network models, we use the $s=10$ preceding days and set $\tilde{\mathbf{x}}_{t}=\{(\mathbf{x}_{j},y_{j})\}_{j=t-s}^{t-1}$. The augmented training set is then $\tilde{\mathcal{D}}=\{(\tilde{\mathbf{x}}_{t},y_{t})\}_{t=s+1}^{T}$. Only observations of the $p+1=8$ variables during the preceding ten days are used to predict one day ahead for a new test time point.
@@ 5.3 Results
The middle panel of Figure 7 shows the validation loss (solid line) of the selected EQRN model as a function of the training epoch. It can be seen that already after a few epochs, the method has a lower loss than the simple semi-conditional model with constant GPD parameters $\sigma$ and $\xi$ (dashed line). This shows that the GPD distribution varies with the predictor values $\tilde{\mathbf{x}}_{t}$ and that a flexible model is beneficial.
@@
Consequently, plugging in the data $(\tilde{\mathbf{x}}_{t},y_{t})_{t=1}^{T}$ and estimates $\hat{Q}_{\tilde{\mathbf{x}}_{t}}(\tau)$ from a quantile regression method, the equation should approximately hold if the model well calibrated. Such a model assessment plot is shown in the right-hand panel of Figure 7 for our EQRN fit as a function of the quantile level $\tau$. We observe that the model is fairly well calibrated, with a slight bias towards more exceedances than expected.
@@
An additional output are the corresponding GPD parameters $\sigma(\tilde{\mathbf{x}}_{u})$ and $\xi(\tilde{\mathbf{x}}_{u})$, which together with the intermediate quantile $\hat{Q}_{\tilde{\mathbf{x}}_{u}}(\tau_{0})$ specify the whole tail of the distribution of $Y_{u}\mid\tilde{\mathbf{X}}_{u}=\tilde{\mathbf{x}}_{u}$ according to (5). For a given threshold level of interest $Q$ we can plot the flood risk for the next day as the one-day-ahead forecast of the exceedance probability over $Q$, that is, an estimate of the function $u\mapsto\mathbb{P}(Y_{u}>Q\mid\tilde{\mathbf{X}}_{u}=\tilde{\mathbf{x}}_{u})$. The bottom panel of Figure 2 shows the EQRN-based estimate of this function on the test set as a ratio to the unconditional $\mathbb{P}(Y_{u}>Q)$, where the threshold $Q$ is chosen as the static 100-year return level $Q^{100}$ based on the GEV distribution fitted on the training set; this threshold is relevant since it is often used to determine the height of dams for flood management. It results in a daily measure of how likely the exceedance on the next day is compared to what was expected unconditionally. Times with large predicted probability ratios are apparently times of imminent danger that can be used as triggers for early warning systems or additional flood management measures.
@@
As an example, one may issue a warning when the forecasted conditional probability of exceeding $Q^{100}$ is, say, a hundred times larger than the baseline unconditional probability $\mathbb{P}(Y_{u}>Q^{100})$. In the test data, there are four time clusters, that typically last several days, when $Q^{100}$ is exceeded; see Figure 8 for one of these events. Applying this early warning system, in all four of these cases a timely warning would have been issued on the days preceding the first exceedance of the cluster. Such a decision rule would lead to an average of only $1.3$ warnings for clusters of exceedances per year on the test set and is therefore not overly conservative. As the model does not need re-fitting on the test set, the daily forecast and possible warnings are obtained in less than one second of computation time, even on a CPU-only laptop computer [^5] Intel Core i5-8265U 1.6 GHz processor with 4 cores, and 8 GB of RAM memory.. The training time for the selected GPD network took less than 15 minutes.
@@
We consider the period of the 2005 flood in Bern in more detail. Figure 8 shows the two discharge time series and precipitation at the closest meteorological station before and after the event. On the evening of August 21, 2005, the day preceding the first exceedance of the 100-year GEV return level (horizontal dashed line), the prediction of our EQRN already indicated a sudden increase in the probability of this exceedance; see bottom panel of Figure 2. Equivalently, the blue point in Figure 8 shows the increased value of the conditional 100-year quantile predicted by the model. The diamonds mark the observations of the previous ten days that were used for this prediction.
@@
In this case, the high precipitation values on August 21 and the preceding days seem to have driven this prediction, possibly together with high values of the rivers. It is interesting to note that a similar situation on August 2 has not resulted in a “flood warning” since the forecasted exceedance probability and return level are not exceptionally high – as a matter of fact, there was no exceedance on the next day. This means that the predictions are driven by a complex combination of the risk factors $\tilde{\mathbf{X}}_{t}$ in space and time that are well captured by the recurrent architecture of the EQRN.
@@
At the time, the FOEN used an adapted version of the hydrological model HBV (Lindström et al., 1997) for forecasting river discharges based on several inputs such as precipitation forecasts. However, the forecasts prior to this event underestimated the flood risk and resulted in too-late warnings. Physical models for discharges and precipitation do not use explicit extrapolation in the extreme tails and often have poor performance in the largest predictions. In fact, during the 2005 flood, a main reason for the late warnings was that forecasters did not trust the predicted precipitation amounts during this extreme scenario. In the aftermath of this flood, the FOEN therefore published a detailed analysis of the internal forecasting procedures (Bezzola and Hegg, 2007). The exact forecasts from that time are not available, but the above discussion of the results shows that our statistical methodology is a competitive alternative to physical models for the forecasting of flood risk. We also note that our model uses much less information as it relies only on observed discharge and precipitation and does not require forecasts of atmospheric variables.
In Section S.5 of the Supplementary Material, we compare and discuss the forecasts of our EQRN method with those of some of the competitors. Figures S.8-S.11 show that all methods seem to capture at least some of the temporal structure based on the past covariates. Except for the GBEX method, the forecasts of the competitors suffer however from a low sensitivity to changes in the conditional tail or from a too erratic behaviour as a function of time. Overall, the recurrent structure of our EQRN method therefore seems to be the best suited model for this kind of sequentially dependent data. Since in real-world applications the true quantiles are unknown, direct computation of the prediction error is difficult and model assessment plots as in the right-hand panel of Figure 7 are crucial. This highlights the importance of simulation studies to evaluate and compare quantitatively the accuracy of different methods. In particular in situations with temporal dependence, our EQRN method clearly outperforms the competitors (e.g., Figure 5). This is another indicator to trust the EQRN forecasts in applications.

### v4

@@ 4.2 Results
For estimation of the conditional GPD parameters with our recurrent EQRN model, we consider both LSTM and GRU architectures with one to three recurrent layers and hidden dimensions between $32$ and $256$. As the networks are not too deep, $L_{2}$ penalty was chosen over dropout for regularization during training, with possible penalty $\lambda\in\{0,10^{-6},10^{-5},10^{-4},10^{-3}\}$. Both constant and covariate-dependent shape parameter outputs are considered. The model with minimum validation loss is a single LSTM layer with hidden dimension $128$, followed by the usual fully connected output layer, with constant shape and $\lambda=10^{-4}$. As a comparison we also retain predictions for the best unpenalized network with $\lambda=0$ and fixed shape, which has two LSTM layers of hidden size $128$.
The left panel of Figure 5 shows the RMSE of best penalized and unpenalized EQRN models, compared to the improved competitors, as a function of the quantile level $\tau$. We observe that, for the lowest level $\tau=\tau_{0}=80\%$, all structured GPD models have the same performance since they use the same intermediate quantiles. GRF and the unconditional model have already higher errors since they are not able to capture the sequential dependence at the intermediate level sufficiently. For growing quantile levels $\tau$, the errors of the covariate-dependent GPD models start to diverge. This is due to the differences in modelling flexibility in terms of the GPD parameters of each method. We observe a similar behaviour for EXQAR, as its linear quantile dependence is not flexible enough. Our EQRN method based on recurrent neural networks seems to be best at modelling sequential tail dependence.
Figure 5: Left: Root mean squared error between predicted and true conditional quantiles at different probability levels $\tau$ (log-scale), for the selected EQRN models and the improved competitors. Centre-right: True vs. predicted quantiles at probability level $\tau=99.95\%$ for the best unpenalized (middle) and penalized (right) EQRN models (dots), compared to the semiconditional estimates (crosses).
Figure 5 also shows the predicted quantiles $\hat{Q}_{\tilde{\mathbf{x}}_{u}}(\tau)$ on the test data compared to the true $Q_{\tilde{\mathbf{x}}_{u}}(\tau)$ for a fixed $\tau=99.95\%$ for the best penalized and unpenalized EQRN models. In general, both models seem to perform well in predicting the high conditional quantiles. The weight penalty seems to mainly affect the larger quantile predictions. Compared to the unpenalized model, we observe that the reduction in the variance of the predictions comes at the cost of a bias for larger quantile values. This bias-variance trade-off is typical with penalization. The poor performance of the semiconditional estimates highlights the added value of covariate dependence in the GPD parameters.
@@ 5.1 Motivation
Flood risk is a major natural hazard in Europe, which causes huge economic damage and endangers human lives. There is a longstanding interest in statistical methods of extreme value theory for hydrology (e.g., Katz et al., 2002; Keef et al., 2009; Asadi et al., 2015; Engelke and Hitz, 2020), and national agencies commonly use them to assess the long-term risk of flooding in cities, at power plants, and other key locations. Return levels with long return periods can be estimated using the GEV distribution for annual maxima or the GPD model for daily threshold exceedances. The output then guides effective long-term flood management measures.
An example of an important location in Switzerland is the gauging station in Bern on the Aare river, which is shown within its water catchment in Figure 1. The Swiss Federal Office for the Environment (FOEN) monitors the Aare, and we use daily average discharges (in $\text{m}^{3}\text{s}^{-1}$) in Bern and another upstream station together with recordings of daily precipitation (in mm) at six locations in the Bern catchment; see Figure 1 for details. All time series are available in the period from 1930–2014 and can be obtained from the FOEN [^2] https://www.hydrodaten.admin.ch/. (for discharges) and MeteoSwiss [^3] https://gate.meteoswiss.ch/idaweb. (for precipitation). Figure 6 shows an excerpt for the two river stations and one precipitation gauge.
To illustrate possible drawbacks of a classical extreme value analysis, the left panel of Figure 7 shows the annual maxima of river discharges at the Bernese station on the Aare. The dashed line is the estimated 100-year return level based on the GEV approximation using the training period from 1930–1958. One can see that starting from the year 1999 there are several exceedances over this return level, somewhat contradicting the fact that it should only be exceeded on average once in 100 years. The solid line is the same return level based on data from 1930–$y$, where $y\in\{1959,\dots,2014\}$ denotes the end of the training period. While the predictions are fairly stable until 1999, an extreme value analysis performed after that year would yield much higher values for the 100-year return level. Conversely, historical estimates before such a break-point would severely underestimate the flood risk. In general, distributional shifts can be due to climate change, changes in the river system, or other structural breaks in factors influencing discharge at this location. For the Bernese station, the FOEN indeed reports a significant break-point in extreme discharges in the nineties but acknowledges that a clear cause can not be identified [^4] See flood report of the FOEN at https://www.hydrodaten.admin.ch/en/2135.html.. One factor may be a multidecadal variability of flood occurrence, as described in Schmocker-Fackel and Naef (2010).
@@
Figure 7: Left: Annual maxima of daily average discharges (points) at Bern–Schönau (62) together with the unconditional 100-year return level based on GEV fitted on the training data of 1930–1958 (dashed line), and the evolution of the same return level (solid line) using data from 1930–$y$, where $y\in\{1959,\dots,2014\}$ denotes end of the period. Middle: Evolution of the validation loss (solid line) of the selected EQRN network for the river discharge data as a function of the training epoch; the dashed line shows the validation loss of the semiconditional model. Right: Number of observations exceeding the EQRN quantile predictions on the test set (dotted line) compared to the expected number of exceedances (solid line) for different probability levels (log-scale).
The aim of our methodology is complementary to classical extreme value analysis and addresses this issue with static return levels. We apply our EQRN model to estimate one-day-ahead extreme quantiles of the river discharge conditionally on previous observations of discharge and precipitation in the catchment. This allows the forecasting of flood risk even in nonstationary systems, such as a changing climate. The strength of our approach lies in the ability to exploit information from multiple covariates and capture the complex time dependence. Even in situations where the causes for structural changes are unknown, our method implicitly accounts for them through their effects on the covariates. The output of the model can help practitioners and authorities to manage flooding more effectively and help to minimize their disastrous impacts by early warning systems.
@@ 5.2 Model specification
The whole dataset consists of 31,046 daily observations $(\mathbf{x}_{t},y_{t})$ between 1930–2014. The response $y_{t}$ is the daily average discharge at the Bernese gauging station on the Aare, and the covariates $\mathbf{x}_{t}\in\mathbb{R}^{p}$, $p=7$, consist of discharge at another upstream station and daily precipitation measurements from six locations in the same catchment; see Figures 1 and 6. The discharges show significant seasonality, both in trend and variance, with the largest extremes only appearing in the summer. We do not reduce artificially the nonstationarity via classical approaches from times series analysis (e.g., Cleveland et al., 1990), as we believe the seasonality and other trends are captured through the covariates.
We split the data into training and test sets. The first $T=10,349$ observations in the period between 1930–1958 are used to train the models, whereof the first three-quarters serve the parameter estimation, and the remaining quarter is a validation set to determine hyperparameters (sets $\mathcal{T}$ and $\mathcal{V}$ in Algorithm 2, respectively). The test set contains 20,697 observations from 1958–2014, which is used for neither fitting nor selection of parameters but only to evaluate the model performance on an independent time period. We choose this rather small proportion of training data to study the ability of the model to adapt to possible nonstationarity over time without refitting. In particular, the model weights are not updated with any information from data after 1958, even for forecasts in the study of the 2005 flood of interest. A large test set is also required to evaluate extreme properties of the data distribution. As augmented covariates at time $t$ for the recurrent neural network models, we use the $s=10$ preceding days and set $\tilde{\mathbf{x}}_{t}=\{(\mathbf{x}_{j},y_{j})\}_{j=t-s}^{t-1}$. The augmented training set is then $\tilde{\mathcal{D}}=\{(\tilde{\mathbf{x}}_{t},y_{t})\}_{t=s+1}^{T}$. Only observations of the $p+1=8$ variables during the preceding 10 days are used to predict one day ahead for a new test time point.
@@ 5.3 Results
The middle panel of Figure 7 shows the validation loss (solid line) of the selected EQRN model as a function of the training epoch. It can be seen that already after a few epochs, the method has a lower loss than the simple semiconditional model with constant GPD parameters $\sigma$ and $\xi$ (dashed line). This shows that the GPD distribution varies with the predictor values $\tilde{\mathbf{x}}_{t}$ and that a flexible model is beneficial.
@@
Consequently, plugging in the data $(\tilde{\mathbf{x}}_{t},y_{t})_{t=1}^{T}$ and estimates $\hat{Q}_{\tilde{\mathbf{x}}_{t}}(\tau)$ from a quantile regression method, the equation should approximately hold if the model is well-calibrated. Such a model assessment plot is shown in the right-hand panel of Figure 7 for our EQRN fit as a function of the quantile level $\tau$. We observe that the model is fairly well-calibrated, with a slight bias toward more exceedances than expected.
@@
An additional output are the corresponding GPD parameters $\sigma(\tilde{\mathbf{x}}_{u})$ and $\xi(\tilde{\mathbf{x}}_{u})$, which together with the intermediate quantile $\hat{Q}_{\tilde{\mathbf{x}}_{u}}(\tau_{0})$ specify the whole tail of the distribution of $Y_{u}\mid\tilde{\mathbf{X}}_{u}=\tilde{\mathbf{x}}_{u}$ according to (5). For a given threshold level of interest $Q$, we can plot the flood risk for the next day as the one-day-ahead forecast of the exceedance probability over $Q$, that is, an estimate of the function $u\mapsto\mathbb{P}(Y_{u}>Q\mid\tilde{\mathbf{X}}_{u}=\tilde{\mathbf{x}}_{u})$. The bottom panel of Figure 2 shows the EQRN-based estimate of this function on the test set as a ratio to the unconditional $\mathbb{P}(Y_{u}>Q)$, where the threshold $Q$ is chosen as the static 100-year return level $Q^{100}$ based on the GEV distribution fitted on the training set; this threshold is relevant since it is often used to determine the height of dams for flood management. It results in a daily measure of how likely the exceedance on the next day is compared to what was expected unconditionally. Times with large predicted probability ratios are apparently times of imminent danger that can be used as triggers for early warning systems or additional flood management measures.
@@
As an example, one may issue a warning when the forecasted conditional probability of exceeding $Q^{100}$ is, say, a hundred times larger than the baseline unconditional probability $\mathbb{P}(Y_{u}>Q^{100})$. In the test data, there are four time clusters, which typically last several days, when $Q^{100}$ is exceeded; see Figure 8 for one of these events. Applying this early warning system, in all four of these cases a timely warning would have been issued on the days preceding the first exceedance of the cluster. Such a decision rule would lead to an average of only $1.3$ warnings for clusters of exceedances per year on the test set and is, therefore, not overly conservative. As the model does not need refitting on the test set, the daily forecast and possible warnings are obtained in less than one second of computation time, even on a CPU-only laptop computer [^5] Intel Core i5-8265U 1.6 GHz processor with four cores, and eight gigabytes of RAM memory.. The training time for the selected GPD network took less than 15 minutes.
@@
We consider the period of the 2005 flood in Bern in more detail. Figure 8 shows the two discharge time series and precipitation at the closest meteorological station before and after the event. On the evening of August 21, 2005, the day preceding the first exceedance of the 100-year GEV return level (horizontal dashed line), the prediction of our EQRN already indicated a sudden increase in the probability of this exceedance; see bottom panel of Figure 2. Equivalently, the blue point in Figure 8 shows the increased value of the conditional 100-year quantile predicted by the model. The diamonds mark the observations of the previous 10 days that were used for this prediction.
@@
In this case the high precipitation values on August 21 and the preceding days seem to have driven this prediction, possibly together with high values of the rivers. It is interesting to note that a similar situation on August 2 has not resulted in a “flood warning” since the forecasted exceedance probability and return level are not exceptionally high—as a matter of fact, there was no exceedance on the next day. This means that the predictions are driven by a complex combination of the risk factors $\tilde{\mathbf{X}}_{t}$ in space and time that are well-captured by the recurrent architecture of the EQRN.
@@
At the time the FOEN used an adapted version of the hydrological model HBV (Lindström et al., 1997) for forecasting river discharges based on several inputs such as precipitation forecasts. However, the forecasts prior to this event underestimated the flood risk and resulted in too-late warnings. Physical models for discharges and precipitation do not use explicit extrapolation in the extreme tails and often have poor performance in the largest predictions. In fact, during the 2005 flood, a main reason for the late warnings was that forecasters did not trust the predicted precipitation amounts during this extreme scenario. In the aftermath of this flood, the FOEN, therefore, published a detailed analysis of the internal forecasting procedures (Bezzola and Hegg, 2007). The exact forecasts from that time are not available, but the above discussion of the results shows that our statistical methodology is a competitive alternative to physical models for the forecasting of flood risk. We also note that our model uses much less information, as it relies only on observed discharge and precipitation and does not require forecasts of atmospheric variables.
In Section S.5 of the Supplementary Material, we compare and discuss the forecasts of our EQRN method with those of some of the competitors. Figures S.8S.11 show that all methods seem to capture at least some of the temporal structure based on the past covariates. Except for the GBEX method, the forecasts of the competitors suffer, however, from a low sensitivity to changes in the conditional tail or from a too erratic behaviour as a function of time. Overall, the recurrent structure of our EQRN method, therefore, seems to be the best suited model for this kind of sequentially dependent data. Since in real-world applications the true quantiles are unknown, direct computation of the prediction error is difficult, and model assessment plots as in the right-hand panel of Figure 7 are crucial. This highlights the importance of simulation studies to evaluate and compare quantitatively the accuracy of different methods. In particular, in situations with temporal dependence, our EQRN method clearly outperforms the competitors (e.g., Figure 5). This is another indicator to trust the EQRN forecasts in applications.